# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ III Thermodynamic performance
In any of the three cases, the TUR implies that beyond a certain threshold, a classical Markovian engine can only enhance its precision in the output (cooling) power at the cost of either reducing the output itself or reducing the energy efficiency, so that the above ratio remains bounded by $2$, i.e. $\mathcal{Q}\geq 2$. However, some models of quantum thermal machines have been shown to produce violations of the TUR, that is, they verify $\mathcal{Q}<2$ (see e.g. Refs. [28, 29, 30, 31, 32, 33]). Such violations act as a witness indicating an enhanced tradeoff between power, precision and efficiency that arise in certain parameter regimes. Nevertheless, it is in principle unclear whether that violations may occur in relevant regimes, i.e. where the machine performs in an optimal way, providing a practical quantum-thermodynamic advantage. In addition, such quantum-thermodynamic advantages may arise even if the TUR is not violated, since classical machines may not saturate the TUR in such relevant regimes. Hence in order to provide a fair and accurate assessment of quantum-thermodynamic advantages in thermal machines both optimization of the machine performance and comparison to classical models become necessary.
IV Classical thermodynamic equivalents of thermal machines
Throughout this paper, inspired by the notion of classical emulability introduced in Ref. [23], we use the term “classical thermodynamic-equivalent” (or simply “classical equivalent”) of a quantum thermal machine to refer to a thermal machine model with same bare Hamiltonian, $H_{0}$, but whose evolution can be described using only classical Markovian dynamics (stochastic jumps between the energy levels), while being capable of producing the same average currents as the quantum machine, by using the same amount of incoherent resources. Notice that if the classical equivalent machine is capable of producing the same average currents, this ensures the same power output and efficiency as the original quantum machine. However, quantum and classical equivalent models may different in general in their fluctuations.
By using same incoherent resources we mean that the classical-thermodynamic equivalent model is in contact with the same thermal baths, and therefore have access to the same temperatures. Furthermore, the requisite of having same bare Hamiltonian $H_{0}$ implies that the classical equivalent also has the same energy level structure as the original quantum machine, which allow us to identify them as the “same machine”. This is in contrast with other notions of classical analogs in thermal machines where all the parameters of the classical model are imposed to be exactly equal, including the couplings to the baths or the driving strength [9, 12] (this is also the case of introducing extra dephasing in the quantum model). Here instead we allow to vary these parameters as long as the requirements for weak-coupling and Markovian dynamics assumed for the thermal machine models are satisfied. This choice is not arbitrary, but allows for an stringer notion of genuine quantum-thermodynamic advantage that avoids spurious “advantages” that may disappear by just slightly modifying some of the parameters in the classical model.
In the following, we develop a general method for constructing these equivalent machines in situations where the coherence in the system arises either from Hamiltonian dynamics (Hamiltonian-induced coherence) or from dissipative processes (noise-induced coherence).
IV.1 Hamiltonian-induced coherence
Since we want the classical equivalent to produce the same average energy currents than the quantum counterpart, our starting point will be the generic expression for the heat currents given in Eq. (13). This expression, valid for any given (bare) Hamiltonian $H_{0}$, can be rewritten as (see App. A):
where we remind the reader that transitions in the sum above are restricted to the ones induced in the set $B_{r}$, i.e. induced by reservoir $r$. The above expression only depends on the level populations (diagonal elements of the density matrix), energy gaps of the machine and jump rates. As a consequence, the useful output current of the machine can only depend on these quantities as follows from the first law, e.g. in heat engines $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$. Given that the energy gaps and temperatures of the reservoirs are fixed, we conclude that to ensure identical input and output currents, the quantum machine and its classical equivalent must have matching diagonal elements in their steady-state density matrices.
In order to mimic the same level populations in the classical equivalent, we proceed by first solving the equations of motion for the diagonal elements and non-vanishing coherences in the quantum machine (details of this procedure are presented in Appendix C). Let’s denote by indices |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$ a couple of levels that are connected by $V$ (for simplicity we assume |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$ not further connected to other levels). The equations of motion for these two levels read:
$\begin{split}\dot{\rho}_{\mathrm{uu}}&=\sum_{j\neq\mathrm{u}}\gamma_{j\mathrm{u}}\rho_{jj}-\rho_{\mathrm{uu}}\sum_{i}\gamma_{\mathrm{u}i}-2g\,\text{Im}\left(\rho_{\mathrm{uv}}\right),\\ \dot{\rho}_{\mathrm{vv}}&=\sum_{j\neq\mathrm{v}}\gamma_{j\mathrm{v}}\rho_{jj}-\rho_{\mathrm{vv}}\sum_{i}\gamma_{\mathrm{v}i}+2g\,\text{Im}\left(\rho_{\mathrm{uv}}\right),\\ \dot{\rho}_{\mathrm{uv}}&=-\dfrac{1}{2}\sum_{i}\left(\gamma_{\mathrm{u}i}+\gamma_{\mathrm{v}i}\right)\rho_{\mathrm{uv}}-ig\left(\rho_{\mathrm{vv}}-\rho_{\mathrm{uu}}\right),\end{split}$ (23)
Following Ref. [23], by equating all the derivatives to zero in Eqs. (23), we can determine the relation between the coherence of levels connected by the Hamiltonian $V$ and their populations (that should be verified in the steady state):
$\pi_{\mathrm{uv}}=\dfrac{-2ig\left(\pi_{\mathrm{vv}}-\pi_{\mathrm{uu}}\right)}{\sum_{i}\left(\gamma_{\mathrm{u}i}+\gamma_{\mathrm{v}i}\right)}.$ (24)
Then we introduce the above dependence back into off-diagonal elements in Eqs. (23), to obtain that the net effect of coherence in the steady state is equivalent to adding a virtual transition promoting jumps between the interacting levels |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$:
with transition rates
$\gamma_{\mathrm{uv}}^{\rm cl}=\gamma_{\mathrm{vu}}^{\rm cl}=\dfrac{4g^{2}}{\sum_{i}\left(\gamma_{\mathrm{u}i}+\gamma_{\mathrm{v}i}\right)}.$ (26)
As a consequence, if we replace the Hamiltonian term $V$ responsible for the coherent interaction between levels |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$ by the above extra stochastic transition between them, the system governed by Eqs. (25) will reach an steady-state with exactly the same populations as the original one governed by Eq. (23), and therefore the same currents in Eq. (22).
We have thus achieved a classical equivalent of the quantum model in which the dynamics is entirely classical, as given by a set of incoherent jumps between the machine energy levels, while reproducing the same (average) currents in the steady state. We also notice that the equivalent machine is coupled to the same set of thermal baths, while the coherent interaction is replaced by an stochastic jump process without bias in either direction. This provides a classical equivalent model that uses the same amount of incoherent resources. In the case of energetic coherence, i.e. non-degenerate levels |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$ ($\epsilon_{\mathrm{u}}\neq\epsilon_{\mathrm{v}}$) this means replacing the quantum-coherent work source (battery) by a classical stochastic work source (battery), like a bath at infinite temperature. On the other hand, in the case of degenerate coherence ($\epsilon_{\mathrm{u}}=\epsilon_{\mathrm{v}}$) the extra stochastic transition is a source of pure noise without any associated energy current and can hence be considered as free.
IV.2 Noise-induced coherence
@@
$\begin{split}\langle\dot{Q}_{r}\rangle=&\sum_{i<j}(\epsilon_{j}-\epsilon_{i})(\gamma_{ij}\pi_{ii}-\gamma_{ji}\pi_{jj})\\ &+2\sum_{j}(\epsilon_{j}-\epsilon_{\mathrm{v}})(\gamma_{\mathrm{u}j}+\gamma_{\mathrm{v}j})\mathrm{Re}(\pi_{\mathrm{uv}}),\end{split}$ (27)
while the currents that are not involved in the noise-inducing mechanism are given by Eq. (22) as in the previous case (see App. A). Moreover, in Appendix C we show that reproducing the steady-state populations results again in equal average currents, despite of the appearance of the second contribution in Eq. (27).
Following the same procedure as for Hamiltonian-induced coherence, we can observe the net effect of coherence in the steady state by solving the equations of motion and replacing the dependence of coherence back into the equations (see details in App. C):
where $i=\mathrm{u},\mathrm{v}$ above and we defined $\gamma_{\mathrm{uv}}^{*}:=\sum_{j}\sqrt{\gamma_{\mathrm{u}j}\gamma_{\mathrm{v}j}}/\sum_{j}(\gamma_{\mathrm{u}j}+\gamma_{\mathrm{v}j})$. It can be proved that local detailed balance relations are not modified in any transition of the machine by the corrections to the rates above. To show this, let us rewrite the rates of the collective transitions to and from levels $i=\{\mathrm{u},\mathrm{v}$} as $\gamma_{in}=\gamma_{r}^{i}\exp[{\beta_{r}(\epsilon_{i}-\epsilon_{n})/2}]$ and $\gamma_{ni}=\gamma_{r}^{i}~{}\exp[{-\beta_{r}(\epsilon_{i}-\epsilon_{n})/2}]$ respectively, were $\gamma_{r}^{i}:=\sqrt{\gamma_{in}\gamma_{ni}}$ is a purely kinetic (spontaneous-emission-like) contribution to the rates not depending on the direction of the jumps, and $\beta_{r}$ is the temperature of the bath to which the transition is coupled. Using this notation the corrected rates read

### v4

@@ Appendix B Thermodynamic performance
In any of the three cases, the TUR implies that beyond a certain threshold, a classical Markovian engine can only enhance its precision in the output (cooling) power at the cost of either reducing the output itself or reducing the energy efficiency, so that the above ratio remains bounded by $2$, i.e. $\mathcal{Q}\geq 2$. However, some models of quantum thermal machines have been shown to produce violations of the TUR, that is, they verify $\mathcal{Q}<2$ (see e.g. Refs. [28, 29, 30, 31, 32, 33]). Such violations act as a witness indicating an enhanced trade-off between power, precision, and efficiency that arises in certain parameter regimes. Nevertheless, such quantum-thermodynamic advantages may arise even if the TUR is not violated, since classical machines may not saturate the TUR in relevant operational regimes. Hence in order to provide a fair and accurate assessment of quantum-thermodynamic advantages in thermal machines, comparison to classical models becomes necessary.
Appendix C General expression of current first moments with local dissipation
In this Appendix we derive a general expression for the average heat currents including the cases in which coherence may be generated either by Hamiltonian or noise-induced sources. We start by expanding the expression for the standard definition for the heat currents [Eq.  (33)], that is:
$\langle\dot{Q}_{r}\rangle=\sum_{k}\operatorname{Tr}[(H_{0}+V(t))\mathcal{D}_{k}^{(r)}[\pi(t)]]$ (42)
$\simeq\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]]=\sum_{k}\operatorname{Tr}[\mathcal{D}_{k}^{\dagger(r)}[H_{0}]\pi(t)]$
$=\sum_{k}\sum_{m,l}\bra{m}L_{k}^{(r)\dagger}H_{0}L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},H_{0}\}\ket{l}\pi_{lm}$
where the third-order terms $\operatorname{Tr}[V(t)\mathcal{D}_{k}^{(r)}[\pi(t)]]\sim g\gamma_{0}^{(r)}$, with $|V|\sim g$ and $\gamma_{k}^{(r)}$ being the spontaneous emission rate in reservoir $r$ and transition $k$, have been neglected in the second line in accordance to the weak coupling and weak driving approximations. In the second line above, we have applied the cyclic property of the trace to obtain the adjoint dissipator $\mathcal{D}_{k}^{\dagger(r)}[\cdot]:=L_{k}^{(r)\dagger}\cdot L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},\cdot\}$.
Using the explicit form of the Lindblad operators in Eq. (4) and the bare Hamiltonian, $H_{0}=\sum_{i}\epsilon_{i}\ket{i}\bra{i}$ |, we can obtain each term in the sum of the last expression:
$\begin{split}&L_{k}^{(r)\dagger}H_{0}L_{k}^{(r)}=\sum_{i,j,n}\alpha_{ij}^{k}\alpha_{nj}^{k}\sqrt{\gamma_{ij}\gamma_{nj}}\epsilon_{j}\ket{i}\bra{n},\\ &L_{k}^{(r)\dagger}L_{k}^{(r)}H_{0}=\sum_{i,j,n}\alpha_{nj}^{k}\alpha_{ij}^{k}\sqrt{\gamma_{nj}\gamma_{ij}}\epsilon_{i}\ket{i}\bra{n},\\ &H_{0}L_{k}^{(r)\dagger}L_{k}^{(r)}=\sum_{i,j,n}\alpha_{nj}^{k}\alpha_{ij}^{k}\sqrt{\gamma_{nj}\gamma_{ij}}\epsilon_{n}\ket{n}\bra{i},\end{split}$ (43)
where we can use the fact that the Lindblad operators produce jumps only between levels with a fixed energy gap $\Delta\epsilon_{k}=\pm\Delta\epsilon_{r}$, determined by the reservoir, to rewrite the $\alpha$-terms as delta functions:
$\begin{split}&\alpha_{ij}^{k}\alpha_{nj}^{k}=\alpha_{nj}^{k}\alpha_{ij}^{k}=\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{jn}-\Delta\epsilon_{k}),\end{split}$ (44)
where $\Delta\epsilon_{ij}=\epsilon_{i}-\epsilon_{j}$. Replacing Eq. (44) into Eq. (43) we have:
where we used $\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{jn}-\Delta\epsilon_{k})=\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{jn})=\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{ni})$.
By introducing the expressions in (45) into (42) we arrive at a general expression for the heat currents valid for both degenerate and non-degenerate level structures in the machine:
@@ Appendix C General expression of current first moments with local dissipation
$\begin{split}\langle\dot{Q}_{r}\rangle=\sum_{k,n,i,j}^{\in B_{r}}\sqrt{\gamma_{ij}\gamma_{nj}}\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{ni})\\ \times[\epsilon_{j}\pi_{ni}-\epsilon_{n}\mathrm{Re}(\pi_{ni})].\end{split}$ (46)
$\begin{split}\langle\dot{Q}_{r}\rangle=&\sum_{k,i,j}^{\in B_{r}}\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})(\epsilon_{j}-\epsilon_{i})\gamma_{ij}\pi_{ii}\\ &=\sum_{i<j}^{\in B_{r}}(\epsilon_{j}-\epsilon_{i})(\gamma_{ij}\pi_{ii}-\gamma_{ji}\pi_{jj}),\\ \end{split}$ (47)
as given in Eq. (5). On the other hand, for the cases with degenerate energy levels, $\delta(\Delta\epsilon_{ni})$ can be zero even for $n\neq i$, and extra terms are obtained (for more details see the derivation of equations of motion in Appendix D). In the case of a single pair of degenerate levels |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$ we obtain: